Imaging Inc., a developer of radiology equipment, has stock outstanding as follows: 18,000 shares of cumulative preferred 4% stock, $130 par, and 60,000 shares of $5 par common. During its first four years of operations, the following amounts were distributed as dividends: first year, $62,640; second year, $134,560; third year, $161,640; fourth year, $178,800.

Compute the dividends per share on each class of stock for each of the four years. Round all answers to two decimal places. If no dividends are paid in a given year, enter "0".

1st Year 2nd Year 3rd Year 4th Year
Preferred stock (dividend per share) $ $ $ $
Common stock (dividend per share) $ $ $ $
Answer #1

Annual preferred dividend = 18000*130*4% = $93600

1st year 2 3 4
Preferred stock dividend per share $62640 93600+93600-62640 = $124560 $93600 $93600
Common stock dividend per share 0 134560-124560 = $10000 $68040 $85200